a boat takes 2 hours longer to go 45 miles up a river than to return. If the boat's speed in still water is 12 mph, what is the rate of the current.
Can you please show you step this time.

Answer by stanbon!(97)

Downstream Data: distance= 45; rate=12+c; time=45/(12+c)
Upstream Data: distance = 45: rate = 12-c; time= 45/(12-c)
EQUATION:
Time Upstream - Time Downstream = 2 hours
45/(12-c) - 45/(12+c) = 2
Divide thru by 45 to get:
1/(12-c) - 1/(12-c) =2/45
Simplify :
45[(12+c)-(12-c)]= 2[(12-c)(12+c)]
45[2c] =2[144-c^2]
90c = 288-2c^2
c^2+45c-144=0
(c-3)(c+48)=0
current speed = 3 mph
Cheers,
Stan H.